# Consistent text-to-image generation via Scene De-Contextualization

### v1

@@ 6 Conclusion
In this paper, we identify scene contextualization as a key source of ID shift in T2I generation and conduct a formal investigation. Our analysis shows that this contextualization is an inevitable attention-induced phenomenon and the primary driver of ID shift in T2I models. By deriving theoretical bounds on its strength, we provide a foundation for mitigating this effect. Building on these insights, we introduce SDeC, a training-free embedding editing method that suppresses latent scene-ID correlation subspace through eigenvalue stability analysis, yielding refined ID embeddings for more consistent generation. Extensive experiments validate both the effectiveness and generality of the proposed approach. The limitations and future work are elaborated in Appendix-J.
Reproducibility Statement
The code and data will be made available after the publication of this paper.
@@ Appendix A Proof of Theorem 1
Let $\mathcal{H}_{\text{id}},\mathcal{H}_{\text{sc}}\subset\mathbb{R}^{d}$ be the identity/scene semantic subspaces, which have ideal semantic separation: $\mathcal{H}_{\mathrm{id}}\cap\mathcal{H}_{\mathrm{sc}}=\varnothing$. Let $\mathcal{Z}_{\text{id}}\subseteq\mathcal{H}_{\text{id}}$ and $\mathcal{Z}_{\text{sc}}^{k}\subseteq\mathcal{H}_{\text{sc}}$ be the prompt-embedding matrix of $\mathcal{P}_{\text{id}}$ and $\mathcal{P}^{k}_{\text{sc}}$, respectively; the prompt-embedding matrix be partitioned by semantics $\mathcal{Z}=\begin{bmatrix}\mathcal{Z}_{\mathrm{id}};\mathcal{Z}_{\mathrm{sc}}^{k}\end{bmatrix}\in\mathbb{R}^{n\times d}$. For any query $q_{\text{id}}$ associated with the identity, its attention output formulated by Eq. (1) can be specified to
@@
$T_{\rm{sc}}=\Pi_{\mathrm{id}}\big[\alpha_{\mathrm{sc}}^{\top}(\mathcal{Z}_{\mathrm{sc}}W_{V})\big]=\sum_{j\in\mathrm{scene}}\alpha_{j}\;\Pi_{{\mathrm{id}}}\big[(z^{(s)}_{j})^{\top}W_{V}\big]\neq\mathbf{0}.$ (11)
@@ Appendix B Proof of Corollary 1
$T_{\rm sc}=\Pi_{\mathrm{id}}\!\big[\alpha_{\mathrm{sc}}^{\top}(\mathcal{Z}_{\mathrm{sc}}W_{V})\big]=\sum_{j\in\mathrm{scene}}\alpha_{j}\,\Pi_{\mathrm{id}}\!\big(W_{V}z^{(s)}_{j}\big)\neq 0.$ (13)
@@ Appendix D Proof of Corollary 2
$\|\Pi_{\mathrm{id}}W_{V}P_{\cap}\|_{F}^{2}=\|UU^{\top}W_{V}P_{\cap}\|_{F}^{2}~\text{and}~\|W_{V}^{\top}P_{\cap}^{\perp}\Pi_{\mathrm{id}}\|_{F}^{2}=\|W_{V}^{\top}P_{\cap}^{\perp}UU^{\top}\|_{F}^{2}.$ (20)
@@
$\begin{split}\|UU^{\top}W_{V}P_{\cap}\|_{F}^{2}=\mathrm{tr}\!\left(P_{\cap}^{\top}W_{V}^{\top}UU^{\top}W_{V}P_{\cap}\right)=\mathrm{tr}\!\left(U^{\top}W_{V}P_{\cap}P_{\cap}^{\top}W_{V}^{\top}U\right).\end{split}$ (22)
@@
$\begin{split}\|UU^{\top}W_{V}P_{\cap}\|_{F}^{2}=\mathrm{tr}\!\left((U^{\top}W_{V}P_{\cap})(U^{\top}W_{V}P_{\cap})^{\top}\right)=\|U^{\top}W_{V}P_{\cap}\|_{F}^{2}.\end{split}$ (23)
@@ E.1 Parameters setting
SDeC involves three parameters: Trade-off parameter $\beta$ and switching constant $M$ in Eq. (6) and weighting strength $\Omega$ in Eq. (8). We set $(\beta,\,M,\,\Omega)=(10,\,20,\,1)$ cross all experiments.
@@ E.2 Model setting
We achieve ID-preserving image generation by editing the ID prompt embeddings during the inference phase. There is no extra training or optimization imposed on the generative models. In practice, we adopt Stable Diffusion (SD) V1.5111https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 as the backbone model of BLIP-Diffusion, while the pre-trained Stable Diffusion XL (SDXL) [^2] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 is selected as the backbone model for our SDeC and the rest of the comparison approaches.
@@ Appendix F Demonstration of Scene-level Interference in 1Prompt1Story
Appendix G Further Model Analysis
G.1 Qualitative comparison of ablation study
As a supplement to Tab. 3, in Fig. 7, we present a qualitative comparison between methods SDeC w/o soft-estimation, SDeC w/o abs-excursion, and SDeC. It is seen that the qualitative results are consistent with the data in Tab. 3.
In particular, method SDeC w/o abs-excursion performs worse than method SDeC w/o soft-estimation, which supports our previous discussion. Specifically, SDeC w/o abs-excursion disregards the inherent importance of different SVD eigen directions, resulting in greater semantic loss and severe subject deformation (for example, see image #2 in Fig. 7-Right). In contrast, SDeC w/o soft-estimation, which employs a matrix transformation approach, just suffers from less accurate in identifying latent overlapping subspaces, but remain the key semantic information. Thus, it achieves better ID preservation than SDeC w/o abs-excursion.
@@ G.1 Qualitative comparison of ablation study
G.2 Parameter sensitivity analysis
Our SDeC method involves three parameters, including weighting strength $\Omega$ in Eq. (8) and trade-off parameter $\beta$ and switching constant $M$ in Eq. (6). In Fig. 8, we provide the varying curves of DreamSim-F and CLIP-T scores as the three parameters change. As shown in Fig. 8-Left, the DreamSim-FV and CLIP-T scores remain near-steady when a wide range of $[5,20]$, indicating our method does not rely on the careful selection of trade-off parameter $\beta$.
@@ G.2 Parameter sensitivity analysis
Appendix H Supplemental Experiments
H.1 Integrating with other generative tasks
@@
Figure 11: Comparison of combining SDeC with the base models. From top to bottom, there are results of SDXL group, PlayGround-v2.5-1024px-Aesthetic group, RealVisXL-V4.0 group, and Juggernaut-X-V10 group, respectively.
Figure 12: Supplementary qualitative comparison results.
H.2 Consistent story generation with multiple subjects
@@
H.3 Generality to base generative model
Extensive experiments with the SDXL base model show that our SDeC enables semantically meaningful editing. However, if its effectiveness were confined to only a subset of generative models, its practical applicability would be greatly limited. To demonstrate its generality, we integrate SDeC with four representative base models and compare performance before and after integration. Specifically, in addition to SDXL, the other base models include PlayGround-v2.5-1024px-Aesthetic [^3] https://huggingface.co/playgroundai/playground-v2.5-1024px-aesthetic, RealVisXL-V4.0444https://huggingface.co/SG161222/RealVisXL_V4.0, and Juggernaut-X-V10555https://huggingface.co/RunDiffusion/Juggernaut-X-v10.
@@ H.3 Generality to base generative model
Appendix I More qualitative results
As a supplement to Fig. 4, we provide more qualitative comparison results in Fig. 12.
Appendix J Limitation and Future Work
@@ Appendix I More qualitative results
Appendix K An Empirical Study of Nano Banana’s ID-preservation
K.1 Building baseline
As the beginning of our exploration, this baseline experiment generates 12 images using Nano Banana [^6] https://aistudio.google.com/models/gemini-2-5-flash-image under the setting formulated in the Problem Statement of Sec. 3. In practice, we start a new session and sequentially input the 12 prompts into Nano Banana via its official interface in dialogue mode. We set the ID prompt to “A dreamy illustration of a beautiful princess with a kind smile”, and the scene prompts are generated using ChatGPT-5.0777https://chatgpt.com/.
@@
K.2 Validating leveraging prior images as context
@@
[…]
K.3 Comparing with our method

### v2

@@ 6 Conclusion
In this paper, we identify scene contextualization as a key source of ID shift in T2I generation and conduct a formal investigation. Our analysis shows that this contextualization is an inevitable attention-induced phenomenon and the primary driver of ID shift in T2I models. By deriving theoretical bounds on its strength, we provide a foundation for mitigating this effect. Building on these insights, we introduce SDeC, a training-free embedding editing method that suppresses latent scene-ID correlation subspace through eigenvalue stability analysis, yielding refined ID embeddings for more consistent generation. Extensive experiments validate both the effectiveness and generality of the proposed approach. The limitations and future work are elaborated in Appendix-K.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62476169), the Sichuan Science and Technology Program (2025YFMS0024), and the UKRI-AHRC CoSTAR National Lab for Creative Industries Research and Development (AH/Y001060/1).
@@ Appendix A Proof of Theorem 1
Let $\mathcal{H}_{\text{id}},\mathcal{H}_{\text{sc}}\subset\mathbb{R}^{d}$ be the identity/scene semantic subspaces, which have ideal semantic separation: $\mathcal{H}_{\mathrm{id}}\cap\mathcal{H}_{\mathrm{sc}}=\varnothing$. Let $\mathcal{Z}_{\text{id}}\subseteq\mathcal{H}_{\text{id}}$ and $\mathcal{Z}_{\text{sc}}^{k}\subseteq\mathcal{H}_{\text{sc}}$ be the prompt-embedding matrix of $\mathcal{P}_{\text{id}}$ and $\mathcal{P}^{k}_{\text{sc}}$, respectively; the prompt-embedding matrix be partitioned by semantics $\mathcal{Z}=\begin{bmatrix}\mathcal{Z}_{\mathrm{id}};\mathcal{Z}_{\mathrm{sc}}^{k}\end{bmatrix}\in\mathbb{R}^{n\times d}$. For any query $q_{\text{id}}$ associated with the identity, its attention output can be specified to
@@
$\begin{split}T_{\rm{sc}}&=\Pi_{\mathrm{id}}\big[\alpha_{\mathrm{sc}}^{\top}(\mathcal{Z}_{\mathrm{sc}}W_{V})\big]\\ &=\sum_{j\in\mathrm{scene}}\alpha_{j}\;\Pi_{{\mathrm{id}}}\big[(z^{(s)}_{j})^{\top}W_{V}\big]\neq\mathbf{0}.\end{split}$ (11)
@@ Appendix B Proof of Corollary 1
$\begin{split}T_{\rm sc}&=\Pi_{\mathrm{id}}\!\big[\alpha_{\mathrm{sc}}^{\top}(\mathcal{Z}_{\mathrm{sc}}W_{V})\big]\\ &=\sum_{j\in\mathrm{scene}}\alpha_{j}\,\Pi_{\mathrm{id}}\!\big(W_{V}z^{(s)}_{j}\big)\neq 0.\end{split}$ (13)
@@ Appendix D Proof of Corollary 2
$\begin{split}\|\Pi_{\mathrm{id}}W_{V}P_{\cap}\|_{F}^{2}&=\|UU^{\top}W_{V}P_{\cap}\|_{F}^{2}\\ \|W_{V}^{\top}P_{\cap}^{\perp}\Pi_{\mathrm{id}}\|_{F}^{2}&=\|W_{V}^{\top}P_{\cap}^{\perp}UU^{\top}\|_{F}^{2}.\end{split}$ (20)
@@
$\begin{split}\|UU^{\top}W_{V}P_{\cap}\|_{F}^{2}&=\mathrm{tr}\!\left(P_{\cap}^{\top}W_{V}^{\top}UU^{\top}W_{V}P_{\cap}\right)\\ &=\mathrm{tr}\!\left(U^{\top}W_{V}P_{\cap}P_{\cap}^{\top}W_{V}^{\top}U\right).\end{split}$ (22)
@@
$\begin{split}\|UU^{\top}W_{V}P_{\cap}\|_{F}^{2}&=\mathrm{tr}\!\left((U^{\top}W_{V}P_{\cap})(U^{\top}W_{V}P_{\cap})^{\top}\right)\\ &=\|U^{\top}W_{V}P_{\cap}\|_{F}^{2}.\end{split}$ (23)
@@ E.1 Parameters setting
SDeC involves three parameters: Trade-off parameter $\beta$ and switching constant $M$ in Eq. (6) and weighting strength $\Omega$ in Eq. (8). We fix $(\beta,\,M)=(10,\,20)$ in all experiments. For UNet-based (Ronneberger et al., 2015) generative models, we set $\Omega=1$, whereas MMDiT-based (Esser et al., 2024) models use a larger value, $\Omega=10$. A detailed description of the parameter-setting guidelines is provided in Parameter Analysis section (Appendix H.2).
@@ E.2 Model setting
We achieve ID-preserving image generation by editing the ID prompt embeddings during the inference phase. There is no extra training or optimization imposed on the generative models. In practice, we adopt Stable Diffusion (SD) V1.5222https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 as the backbone model of BLIP-Diffusion, while the pre-trained Stable Diffusion XL (SDXL) [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 is selected as the backbone model for our SDeC and the rest of the comparison approaches.
@@ E.3 Implementation details of method SDeC w/o soft-estimation
Figure 8: Case one of supplementary qualitative comparison results.
@@ Appendix F Demonstration of Scene-level Interference in 1Prompt1Story
Appendix G More qualitative results
As a supplement to Fig. 4, we provide more qualitative comparison results in Fig. 8 and Fig. 9. Although without the story-wide context, the images generated by our method in a “one prompt per scene” manner present better ID preservation while well scene matching.
Figure 9: Case two of supplementary qualitative comparison results.
Appendix H Further Model Analysis
H.1 Discussion of scalability
Although the SVD step in SDeC has a theoretical complexity of $O(d^{3})$ ($d$ is dimension of matrix), it does not form a practical bottleneck during inference. In practice, SDeC incurs only negligible overhead, as the generative model itself accounts for more than 90% of the total computation time.
To further improve scalability, particularly for extremely long prompts, several extensions merit investigation in future work. First, approximate SVD techniques, such as Randomized SVD (Halko et al., 2011) with $O(d^{2}\log k)$ complexity or the Nyström approximation (Williams & Seeger, 2001) with $O(dk^{2})$ ($k$ is a small constant), could significantly reduce computational cost while preserving the essential spectral structure. Second, a divide-and-conquer strategy, in which a long prompt is partitioned into $K$ sub-prompts and processed sequentially, would allow the overhead to grow approximately linearly with $K$ (about $K\times 0.61$ seconds). These observations suggest that SDeC remains scalable and practical even for large-scale or dynamically evolving prompting scenarios.
@@ H.1 Discussion of scalability
H.2 Parameter analysis
Our SDeC method involves three parameters, including weighting strength $\Omega$ in Eq. (8) and trade-off parameter $\beta$ and switching constant $M$ in Eq. (6). In this part, we first specify the setting of these hyperparameters for optimal performance, and then conduct a sensitivity analysis.
Guidelines for parameter setting
In our SDeC method, the setting of hyperparameters for optimal performance should follow the guidelines below:
(1) Parameter $\beta_{M}$ is associated with the "forward-and-backward" optimization. Practically, $M$ should exceed a threshold (20) so that the forward phase can capture as many scene-related directions within the ID embedding as possible. Moreover, $\beta$ should satisfy $\beta\gg 0$, enabling maximal restoration of ID-related information in the backward phase. This "forward-and-backward" optimization aims to identify the latent ID-scene correlation space, independent of and generic with any specific generative model. Consequently, the setting of $\beta$ and $M$ remains consistent across different generative architectures.
(2) Parameter $\Omega$ is tied to the architecture of the underlying generative model, as it represents the strength of de-contextualization and is therefore directly related to the generation process. Our experiments indicate that for UNet-based models (e.g., SDXL, PlayGround, RealVisXL-V4.0, Juggernaut-X-V10), $\Omega$ should be small (typically $\Omega=1$). In contrast, DiT-based models (e.g., SD3, Flux) require a much larger value of $\Omega$ (typically $\Omega=10$).
Sensitivity analysis
In Fig. 10, we provide the varying curves of DreamSim-F and CLIP-T scores as the three parameters change. As shown in Fig. 10-Left, the DreamSim-FV and CLIP-T scores remain near-steady when a wide range of $[5,20]$, indicating our method does not rely on the careful selection of trade-off parameter $\beta$.
@@ H.2 Parameter analysis
H.3 Sensitivity to the ordering of scene and ID prompts
In the evaluation experiments, the prompt is arranged naturally: the ID prompt precedes the scene prompt. To investigate the sensitivity to the ID-scene prompt order, we conduct an additional comparison where the scene and ID prompts are reversed. In this way, we can extend the original SDXL and SDeC (rewritten as SDXL-o and SDeC-o) to variation SDXL-r and SDeC-r, respectively.
The comparison results reported in Tab. 5 indicate that the prompt ordering (ID-Scene vs. Scene-ID) has a non-negligible impact on the metrics, particularly for the baseline model. We hypothesize that tokens placed earlier in the prompt receive more conditioning weight in the cross-attention mechanism, driving the observed sensitivity:
1. Baseline (SDXL) behavior: When the prompt order is reversed (Scene-ID), the baseline model shows worse ID consistency (higher DreamSim-F score) but better scene distinction (higher DreamSim-B score). This suggests the scene tokens, when placed earlier, exert a stronger, negative influence on the subject’s identity, while simultaneously making the scene more distinct.
2. SDeC behavior: When SDeC is applied, the negative effect on the ID from prompt reversing is slightly reduced, which is consistent with our objective of suppressing scene-ID correlation. Importantly, the positive effect on scene distinction is maintained, as SDeC leverages the original scene prompt content.
H.4 Analysis of the "forward-and-backward" optimization
In SDeC, the “forward-and-backward” optimization is central to identifying the latent scene–ID correlation subspace within the ID prompt. To better understand this process, we visualize how the SVD eigenvalues evolve relative to their original values during optimization. For clarity, we show an example with five ID tokens, resulting in five corresponding SVD eigen-directions. We present the visualization results in Fig. 11. There are three observations below.
Table 5: Effect of scene-ID prompt order.
Method DreamSim-F↓ CLIP-I↑ DreamSim-B↑ CLIP-T↑
SDXL-o 0.2778 0.8558 0.3861 0.8865
[…]
SDeC-o 0.2589 0.8655 0.3675 0.8946
[…]
Figure 11: SVD eigenvalues evolving dynamics relative to their original values during the "forward-and-backward" optimization. Here, in the “SVD eigen-direction #,” a smaller value of # indicates a larger eigenvalue for that direction.
First, the evolution of the SVD eigenvalue gap against the original value follows our expectation. In the forward phase (step 0-20), the eigenvalue distinction curves gradually grow to pull the ID embedding to the scene embedding. In the subsequent backward phase (step 20-100), despite some vibration (e.g., the blue one), the curves globally converge to zero. This verifies the recovery of the ID-associated component.
Second, in the final step, the eigenvalue gaps exhibit a clear divergence, which allows us to treat the directions with larger gaps as the basis of the ID–scene shared subspace. Moreover, this divergence is correlated with the specific scene prompt. For example, in the top case, the gap is most pronounced along eigen-direction 2 (red), whereas in the bottom case, the divergence appears along eigen-directions 1 and 2. This also indicates that the “forward-and-backward” optimization imposes a scene-specific de-contextualization.
Third, the eigen-directions with mid-range eigenvalues, such as eigen-directions 1 and 2, exhibit larger final gap values and are thus more readily identified as the basis of the ID–scene shared subspace. This behavior naturally emerges from the “forward-and-backward” optimization: compressing the mid-eigenvalue directions simultaneously preserves the dominant ID-embedding information encoded in the largest-eigenvalue directions while maintaining sufficient compression strength to meaningfully influence the generation process.
Appendix I Supplemental Experiments
I.1 Integrating with other generative tasks
@@
Figure 14: Comparison of combining SDeC with the UNet-based generative models. From top to bottom, there are results of SDXL group, PlayGround-v2.5-1024px-Aesthetic group, RealVisXL-V4.0 group, and Juggernaut-X-V10 group, respectively.
Figure 15: Comparison of combining SDeC with the MMDiT-based generative models. Top: Results of SD3 group. Bottom: Results of Flux group.
I.2 Consistent story generation with multiple subjects
@@
I.3 Generality to base generative model
Extensive experiments with the SDXL base model show that our SDeC enables semantically meaningful editing. However, if its effectiveness were confined to only a subset of generative models, its practical applicability would be greatly limited. To demonstrate its generality, we integrate SDeC with four representative base models and compare performance before and after integration. Specifically, in addition to SDXL, the other base models include PlayGround-v2.5-1024px-Aesthetic [^4] https://huggingface.co/playgroundai/playground-v2.5-1024px-aesthetic, RealVisXL-V4.0555https://huggingface.co/SG161222/RealVisXL_V4.0, and Juggernaut-X-V10666https://huggingface.co/RunDiffusion/Juggernaut-X-v10.
@@ I.3 Generality to base generative model
Table 6: Quantitative results obtained by combining SDeC with four UNet-based generative base models (Top) and two MMDiT-based generative base models (Bottom).
Base-model type Method DreamSim-F↓ CLIP-I↑ DreamSim-B↑ CLIP-T↑
UNet SDXL 0.2778 0.8558 0.3861 0.8865
SDXL+SDeC 0.2589 0.8655 0.3675 0.8946
PlayGround-v2.5 0.2567 0.8680 0.3688 0.8799
PlayGround-v2.5+SDeC 0.2272 0.8832 0.3470 0.8994
RealVisXL-V4.0 0.2616 0.8660 0.4075 0.9007
RealVisXL-V4.0+SDeC 0.2435 0.8665 0.3845 0.8942
Juggernaut-X-V10 0.2748 0.8572 0.4295 0.8974
Juggernaut-X-V10+SDeC 0.2369 0.8781 0.4011 0.9077
MMDiT SD3 0.2875 0.8495 0.4236 0.8596
SD3+SDeC 0.2644 0.8612 0.4140 0.8629
Flux 0.2844 0.8516 0.4238 0.8632
Flux+SDeC 0.2653 0.8623 0.4167 0.8648
The base models above are based on the UNet (Ronneberger et al., 2015) architecture. To verify the generality of our method, we further evaluate our approach using two generative models based on MMDiT (Esser et al., 2024): SD3777https://huggingface.co/stabilityai/stable-diffusion-3-medium and Flux [^8] https://huggingface.co/black-forest-labs/FLUX.1-dev. As shown in Fig. 15, in the SD3 group, the model equipped with our SDeC achieves substantial improvements in appearance consistency, including clothing, hair color, posture, and wing features. In the Flux group, the original Flux model shows an ID shift in gender, whereas incorporating our method significantly mitigates this issue.
For a comprehensive comparison, we list the quantitative results of the methods mentioned above in Tab. 6. In each group, the base model equipped with SDeC achieves stronger ID-related performance (DreamSim-F, CLIP-I) and comparable scenario-level results (DreamSim-B, CLIP-T) relative to the original base models. These results confirm SDeC’s effectiveness and architectural independence.
@@
Appendix J An Empirical Study of Nano Banana’s ID-preservation
J.1 Building baseline
As the beginning of our exploration, this baseline experiment generates 12 images using Nano Banana [^9] https://aistudio.google.com/models/gemini-2-5-flash-image under the setting formulated in the Problem Statement of Sec. 3. In practice, we start a new session and sequentially input the 12 prompts into Nano Banana via its official interface in dialogue mode. We set the ID prompt to “A dreamy illustration of a beautiful princess with a kind smile”, and the scene prompts are generated using ChatGPT-5.0101010https://chatgpt.com/.
@@
J.2 Validating leveraging prior images as context
@@
J.3 Comparing with our method in the way of ID-preservation
@@
Figure 20: Comparison results under scene with strong visual shifts.
J.4 Comparing with our method under scene with strong visual shifts
The de-contextualization studied in this paper focuses on the general effects of a scene on identity (e.g., changes in clothing or posture due to a scene description), which covers normal usage scenarios. If the scene is an extreme case involving strong visual shifts (e.g., intense color washes that fundamentally alter the subject’s appearance, far beyond normal shadow changes), would our method still remain effective?
To address this question, we present the outputs of our SDeC combined with the SDXL base model (denoted by SDXL+SDeC), where the scene prompt includes both normal- and strong-visual shifts. For comparison, we also include results generated by Nano Banana.
From Fig. 20, we observe that the newly released commercial model Nano Banana and SDXL+SDeC fail to capture the direction of the light source, which is crucial for determining shadow formation. This indicates that their ability to interpret and realize the strong visual shifts specified in the scene prompt remains limited.
Appendix K Limitation and Future Work
Additionally, a key motivation of this work is the lack of theoretical justification for ID shift in the T2I literature. We provide the first theoretical interpretation of ID shift through the lens of scene contextualization, and offer a practical solution by addressing the underlying correlation. Nonetheless, this represents only the beginning of a deeper exploration into the fundamental mathematical "balance" between identity and scene. In this paper, we observe two pieces of evidence suggesting the existence of a possible balance. One is the trade-off visualized in Fig. 10 (Middle). The other one is the disentanglement-coherence trade-off reflected in Corollary 2. Developing a systematic theoretical framework for this balance constitutes an important and promising direction for future research.
At present, addressing these extreme visual alterations remains a significant challenge for all current generative models. The results in Fig. 20 highlight the limitations of our method in this regard. Consequently, explicitly incorporating the constraints imposed by extreme visual alterations into the prompt-editing process constitutes an interesting direction for future research.